OpenStudy (mai_mai_312):

HELP PLEASE!!! You and your friend are splitting a restaurant bill and leaving a tip. You pay half of the bill plus $5. Your friend pays half of the bill plus 25% of half the bill. You and your friend pay the same total amount. How much is half the bill? 1) Write an equation. LET X BE HALF OF THE BILL 2) Solve the equation. Show your work. 3) Check your solution. Show your work. 4) State the solution in complete sentences.

OpenStudy (irishboy123):

let the bill be B

OpenStudy (irishboy123):

you pay \(U = \frac{1}{2} B + 5\) your friend pays \(F = \frac{1}{2} B + 0.25 \cdot 0.5 B\) you pay the same amount: \( U = F \) Solve that. And then the question is... ....how much is \(\dfrac{B}{2}\)?